# Adaptive Collaborative Correlation Learning-based Semi-Supervised Multi-Label Feature Selection

### v3

@@ 6.1.3 Implementation Details
Several parameters need to be set in Access-MFS and the other compared methods. For Access-MFS, the objective function in Eq. (8) includes five parameters: $\lambda$, $\theta$, $\mu$, $\alpha$ and $\beta$. The regularization parameters $\alpha$ and $\beta$ are automatically determined while solving $\mathbf{S}$ and $\mathbf{P}$ in Eqs. (21) and (25), respectively. The remaining three parameters $\lambda$, $\theta$ and $\mu$ are tuned using a grid search within the range $\{10^{-3},10^{-2},10^{-1},1,10\}$. For all compared methods, the parameters are tuned in accordance with the original papers to achieve optimal results. For each dataset, following the approach of existing semi-supervised feature selection methods [29, 41], we randomly select $\{10\%,20\%,30\%,40\%,50\%\}$ of the instances as the training set from the entire dataset, using the remainder as the testing set. Additionally, the test data also serve as unlabeled data for the semi-supervised multi-label feature selection methods. As determining the optimal number of selected features is still an open problem [42], we set the range for the number of selected features between 100 and 200 with an increment of 10 across all datasets (except for Parkinson dataset, which contains only 91 features). For Parkinson dataset, we vary the number of selected features from 5 to 55, with an incremental size of 5. Then, Multi-Label K-nearest Neighbor (MLKNN), as proposed in [43], is executed on the selected features. Four commonly used evaluation metrics Average Precision (AP) [1], Macro-F1 (MaF) [1], Ranking Loss (RL) [1] and One Error (OE) [1] are employed to assess the performance of different methods. For the Average Precision and Macro-F1 metrics, a higher value indicates better performance, while for Ranking Loss and One Error, a lower value indicates better performance. All methods are implemented in MATLAB R2022b and executed on a desktop with an Intel Corei9-10900, CPU 2.80 GHz and 64 GB RAM. The experiments are independently conducted five times and the average results are reported for comparison.
@@
ccccccccccccccccc Datasets VirusGO Parkinson Medical Langlog Enron Slashdot Recreation Computer
[…]
Access-MFS .$\mathbf{.8918}$$\mathbf{.5051}$$\mathbf{.8079}$$\mathbf{.4223}$$\mathbf{.8365}$$\mathbf{.1951}$$\mathbf{.2914}$$\mathbf{.0092}$$\mathbf{.6198}$$\mathbf{.0737}$$\mathbf{.5859}$$\mathbf{.1756}$$\mathbf{.5377}$$\mathbf{.1455}$$\mathbf{.6536}$$\mathbf{.1152}$
All-Fea .8621.8621.8621.8621.$4577$$7169.$$2586$$7666.$$1366$$1846.$$0011$$6020.$$0645$$5055.$$0455$$4364.$$0457$6278.6278.6278.6278.0587.0587.0587.0587
SMILE .6798.6798.6798.6798.1685.1685.1685.1685 $7135$$2327$$3985.$$0031$$1753.$$0016$$5692.$$0468$$4080.$$0296$$4293.$$0344$6342.6342.6342.6342.0684.0684.0684.0684
FSSRDM .6955.6955.6955.6955.$2555$$6929.$$1814$$4107.$$0067$$2544.$$0010$$5947.$$0633$$4136.$$0346$$4320.$$0451$6285.6285.6285.6285.0493.0493.0493.0493
SCFS .8639.8639.8639.8639.$4163$$6863.$$2340$$4249.$$0070$$2543.$$0007$$5674.$$0486$$4320.$$0341$$4518.$$0502$6274.6274.6274.6274.0635.0635.0635.0635
LEDA .6798.6798.6798.6798.$1685$$7450.$$3248$$4074.$$0032$$1796.$$0010$$5989.$$0616$$4176.$$0384$$4331.$$0437$6261.6261.6261.6261.0525.0525.0525.0525
LSMR .6901.6901.6901.6901.$2595$$6829.$$1991$$4937.$$0300$$2603.$$0008$$5987.$$0659$$4345.$$0394$$4396.$$0401$6266.6266.6266.6266.0553.0553.0553.0553
SMLFS .8412.8412.8412.8412.$3854$$7572.$$2984$$5318.$$0489$$2733.$$0036$$5946.$$0606$$4117.$$0320$$4250.$$0297$6240.6240.6240.6240.0487.0487.0487.0487
SFAM .8543.8543.8543.8543.$4341$$6951.$$2352$$5976.$$0747$$2279.$$0012$$5877.$$0542$$5037.$$0802$$4327.$$0379$6312.6312.6312.6312.0571.0571.0571.0571
SMDR-IC .7768.7768.7768.7768.$3375$$6220.$$1093$$4114.$$0032$$2486.$$0023$$5410.$$0423$$4076.$$0243$$4235.$$0198$6201.6201.6201.6201.0522.0522.0522.0522
SFS-BLL .8442.8442.8442.8442.$4088$$7530.$$3197$$7384.$$1230$$1693.$$0012$$6006.$$0626$$4154.$$0333$$4369.$$0422$6270.6270.6270.6270.0585.0585.0585.0585
LFFS .7884.7884.7884.7884.$3594$$7168.$$1824$$5139.$$0354$$2538.$$0010$$5947.$$0633$$0439.0439$$4320.$$0451$6285.6285.6285.6285.0493.0493.0493.0493
UAFS-BH .8636.8636.8636.8636.$4273$$6594.$$1684$$4664.$$0233$$2662.$$0040$$5521.$$0452$$4010.$$0301$$4782.$$0805$6319.6319.6319.6319.0604.0604.0604.0604
@@
6.2 Experimental Results and Performance Analysis
In this section, we demonstrate the superior performance of the proposed method Access-MFS compared with other competing methods in terms of AP, MaF, RL and OE. Table 6.1.3 summarizes the experimental results of various methods on eight multi-label datasets based on AP and MaF metrics, with a fixed ratio of labeled instances at 40%, selecting 30 features for the Parkinson dataset and 150 features for the remaining datasets. Due to space constraints, the experimental results regarding RL and OE can be found in Table 1 of the supplementary material. From Table 6.1.3, we can see that the proposed method Access-MFS consistently outperforms other methods across all datasets. As to VirusGO, Parkinson and Recreation datasets, Access-MFS achieves an average improvement of over 10% in both AP and MaF metrics compared to all other methods. On Medical dataset, Access-MFS gains over 32% and 15% average improvement in terms of AP and MaF, respectively. As to Slashdot datasets, Access-MFS outperforms other methods with more than 13% improvement in average AP and MaF. On Computer dataset, Access-MFS obtains over 2.5% average improvement in both AP and MaF. On Langlog and Enron datasets, Access-MFS achieves an average improvement of over 3.6% in AP and still outperforms all other methods in MaF. Besides, Table 1 in the supplementary material also shows that the proposed Access-MFS is still better than other methods on all datasets in terms of RL and OE. Furthermore, compared to the baseline method All-Fea, Access-MFS is superior across all datasets in terms of AP, MaF, RL, and OE, demonstrating the effectiveness of Access-MFS. Access-MFS also shows better performance than the supervised method LEFS on all datasets, indicating the effectiveness of the proposed semi-supervised learning strategy in enhancing feature selection performance.
Due to the difficulty in determining the optimal number of selected features, we also present the performance (AP, MaF, RL and OE) of different methods as the number of selected features varies. Since the page limitation, we only report the experimental results for AP and MaF. The results for RL and OE can be found in Figs. 1 and 2 of the supplementary material. Figs. 2 and 3 respectively show the AP and MaF values for different numbers of selected features, with the labeled instance ratio fixed at 40%. As illustrated in these two figures, the proposed Access-MFS method outperforms other methods in most cases, with the number of selected features varying from 5 to 55 on Parkinson dataset and from 100 to 200 on the other datasets. Furthermore, we also investigate the performance of different methods with varying ratios of labeled instances. Figs. 4 and 5 respectively show the AP and MaF of different methods when the ratio of labeled instances ranges from 10% to 50%. The experimental results for RL and OE with varying ratios of labeled instances can be found in Figs. 3 and 4 of the supplementary material, respectively. From these figures, we can see that our method achieves better performance in most cases. In addition, our method performs better as the ratio of labeled instances increases, indicating that our approach effectively leverages labeled information to guide the collaborative learning of sample and label similarity graphs, ultimately enhancing feature selection performance. The superior performance of the proposed method is attributed to the adaptive collaborative learning of similarity graphs from both the sample and label perspectives, which are embedded into a generalized regression model with an extended uncorrelated constraint and a consistency constraint between predicted and existing labels.
@@ 6.3 Parameter Sensitivity Experiment
Figure 7: Convergence curves of Access-MFS on VirusGo, Parkinson, Enron, and Recreation datasets.
@@ 6.4 Convergence Experiment
The theoretical convergence of the proposed Access-MFS algorithm has been established in Section 5.1. This subsection provides experimental validation of its convergence and analyzes the speed of convergence. Fig. 7 displays the convergence curves of Access-MFS on VirusGo, Parkinson, Enron and Recreation datasets, demonstrating the variation of the objective function value with the number of iterations. As illustrated in Fig. 7, the convergence curves steeply fall within a few iterations and approach stability almost within 10 iterations. The experimental results demonstrate that the proposed method can converge effectively.
TABLE III: Means of AP and MaF for different variants of Access-MFS on eight datasets.
{tabu}
cccccccccc Metircs Methods VirusGO Parkinson Medical Langlog Enron Slashdot Recreation Computer
AP Access-MFS .$\mathbf{.8918}$$\mathbf{.8079}$$\mathbf{.8365}$$\mathbf{.2914}$$\mathbf{.6198}$$\mathbf{.5859}$$\mathbf{.5377}$$\mathbf{.6536}$
Accsess-MFS-I .$8624$$7843$$8009$$2797$$5805$$5453$$5109$6477.6477.6477.6477
Accsess-MFS-II .$8659$$7871$$7967$$2871$$5854$$5521$$5100$6453.6453.6453.6453
Access-MFS-III .$8570$$7568$$7753$$2652$$5399$$4502$$4841$6349.6349.6349.6349
[…]
Accsess-MFS-I .$4328$$3878$$1770$$0048$$0688$$1290$$1199$0890.0890.0890.0890
Accsess-MFS-II .$4515$$4054$$1644$$0027$$0682$$1346$$1187$0956.0956.0956.0956
Access-MFS-III .$4276$$3499$$1545$$0009$$0610$$0878$$0970$0803.0803.0803.0803

### v4

@@ 6.1.3 Implementation Details
Several parameters need to be set in Access-MFS and the other compared methods. For Access-MFS, the regularization parameters $\lambda$, $\theta$ and $\mu$ are tuned using a grid search within the range $\{10^{-3},10^{-2},10^{-1},1,10\}$. For all compared methods, the parameters are tuned using the same grid search strategy described in the original papers to achieve optimal results. For each dataset, following the approach of existing semi-supervised feature selection methods [30, 44], we randomly select $\{10\%,20\%,30\%,40\%,50\%\}$ of the instances as the training set from the entire dataset, using the remainder as the testing set. Additionally, the test data also serve as unlabeled data for the semi-supervised multi-label feature selection methods. As determining the optimal number of selected features is still an open problem [8], we set the range for the number of selected features between 100 and 200 with an increment of 10 across all datasets. Then, Multi-Label K-Nearest Neighbor (MLKNN), as proposed in [45], is executed on the selected features. Four commonly used evaluation metrics Average Precision (AP) [1], Macro-F1 (MaF) [1], Ranking Loss (RL) [1] and One Error (OE) [1] are employed to assess the performance of different methods. For the Average Precision and Macro-F1 metrics, a higher value indicates better performance, while for Ranking Loss and One Error, a lower value indicates better performance. All methods are implemented in MATLAB R2022b and executed on a desktop with an Intel Corei9-10900, CPU 2.80 GHz and 64 GB RAM. The experiments are independently conducted five times and the average results are reported for comparison.
@@
Datasets VirusGO Medical Langlog Enron Slashdot Recreation Computer Science
AP MaF AP MaF AP MaF AP MaF AP MaF AP MaF AP MaF AP MaF
Access-MFS $\mathbf{.8918}$ $\mathbf{.5051}$ $\mathbf{.8365}$ $\mathbf{.1951}$ $\mathbf{.2914}$ $\mathbf{.0092}$ $\mathbf{.6198}$ $\mathbf{.0737}$ $\mathbf{.5859}$ $\mathbf{.1756}$ $\mathbf{.5377}$ $\mathbf{.1455}$ $\mathbf{.6536}$ $\mathbf{.1152}$ $\mathbf{.5076}$ $\mathbf{.0746}$
All-Fea .8621.8621 .4577.4577 .7666.7666 .1366.1366 .1846.1846 .0011.0011 .6020.6020 .0645.0645 .5055.5055 .0455.0455 .4364.4364 .0457.0457 .6278.6278 .0587.0587 .4963.4963 .0649.0649
SMILE .6798.6798 .1685.1685 .3985.3985 .0031.0031 .1753.1753 .0016.0016 .5692.5692 .0468.0468 .4080.4080 .0296.0296 .4293.4293 .0344.0344 .6342.6342 .0684.0684 .4686.4686 .0442.0442
FSSRDM .6955.6955 .2555.2555 .4107.4107 .0067.0067 .2544.2544 .0010.0010 .5947.5947 .0633.0633 .4136.4136 .0346.0346 .4320.4320 .0451.0451 .6285.6285 .0493.0493 .4220.4220 .0128.0128
SCFS .8639.8639 .4163.4163 .4249.4249 .0070.0070 .2543.2543 .0007.0007 .5674.5674 .0486.0486 .4320.4320 .0341.0341 .4518.4518 .0502.0502 .6274.6274 .0635.0635 .4367.4367 .0261.0261
LEDA .6798.6798 .1685.1685 .4074.4074 .0032.0032 .1796.1796 .0010.0010 .5989.5989 .0616.0616 .4176.4176 .0384.0384 .4331.4331 .0437.0437 .6261.6261 .0525.0525 .4301.4301 .0130.0130
LSMR .6901.6901 .2595.2595 .4937.4937 .0300.0300 .2603.2603 .0008.0008 .5987.5987 .0659.0659 .4345.4345 .0394.0394 .4396.4396 .0401.0401 .6266.6266 .0553.0553 .4242.4242 .0142.0142
SMLFS .8412.8412 .3854.3854 .5318.5318 .0489.0489 .2733.2733 .0036.0036 .5946.5946 .0606.0606 .4117.4117 .0320.0320 .4250.4250 .0297.0297 .6240.6240 .0487.0487 .4318.4318 .0201.0201
SFAM .8543.8543 .4341.4341 .5976.5976 .0747.0747 .2279.2279 .0012.0012 .5877.5877 .0542.0542 .5037.5037 .0802.0802 .4327.4327 .0379.0379 .6312.6312 .0571.0571 .4105.4105 .0092.0092
SMDR-IC .7768.7768 .3375.3375 .4114.4114 .0032.0032 .2486.2486 .0023.0023 .5410.5410 .0423.0423 .4076.4076 .0243.0243 .4235.4235 .0198.0198 .6201.6201 .0522.0522 .4009.4009 .0060.0060
SFS-BLL .8442.8442 .4088.4088 .7384.7384 .1230.1230 .1693.1693 .0012.0012 .6006.6006 .0626.0626 .4154.4154 .0333.0333 .4369.4369 .0422.0422 .6270.6270 .0585.0585 .4314.4314 .0138.0138
SMDR-MRC .8475.8475 .4028.4028 .7969.7969 .1873.1873 .2472.2472 .0008.0008 .5980.5980 .0620.0620 .4514.4514 .0878.0878 .5111.5111 .1209.1209 .6371.6371 .0823.0823 .4693.4693 .0516.0516
SMFS .8612.8612 .4823.4823 .7785.7785 .1611.1611 .2635.2635 .0011.0011 .5504.5504 .0477.0477 .5467.5467 .1316.1316 .5117.5117 .1181.1181 .6336.6336 .0680.0680 .4618.4618 .0420.0420
SFGR .8555.8555 .3896.3896 .7727.7727 .1342.1342 .2603.2603 .0041.0041 .6042.6042 .0601.0601 .4346.4346 .0391.0391 .4109.4109 .0190.0190 .6306.6306 .0556.0556 .4245.4245 .0177.0177
LFFS .7884.7884 .3594.3594 .5139.5139 .0354.0354 .2538.2538 .0010.0010 .5947.5947 .0633.0633 .4390.4390 .0439.0439 .4320.4320 .0451.0451 .6285.6285 .0493.0493 .4309.4309 .0159.0159
UAFS-BH .8636.8636 .4273.4273 .4664.4664 .0233.0233 .2662.2662 .0040.0040 .5521.5521 .0452.0452 .4010.4010 .0301.0301 .4782.4782 .0805.0805 .6319.6319 .0604.0604 .4669.4669 .0426.0426
TABLE III: Performance comparison between Access-MFS and other methods on eight datasets in terms of RL and OE.
Datasets VirusGO Medical Langlog Enron Slashdot Recreation Computer Science
RL OE RL OE RL OE RL OE RL OE RL OE RL OE RL OE
Access-MFS $\mathbf{.0711}$ $\mathbf{.1677}$ $\mathbf{.0449}$ $\mathbf{.2007}$ $\mathbf{.1721}$ $\mathbf{.8395}$ $\mathbf{.0968}$ $\mathbf{.3203}$ $\mathbf{.1337}$ $\mathbf{.5354}$ $\mathbf{.1655}$ $\mathbf{.5900}$ $\mathbf{.0829}$ $\mathbf{.4281}$ $\mathbf{.1248}$ $\mathbf{.6127}$
All-Fea .1008.1008 .2065.2065 .0517.0517 .3031.3031 .2050.2050 .9363.9363 .1016.1016 .3365.3365 .1685.1685 .6401.6401 .1986.1986 .7267.7267 .0919.0919 .4507.4507 .1268.1268 .6297.6297
SMILE .2053.2053 .5226.5226 .1416.1416 .7335.7335 .2103.2103 .9416.9416 .1085.1085 .3894.3894 .1985.1985 .7716.7716 .1990.1990 .7405.7405 .0903.0903 .4429.4429 .1367.1367 .6573.6573
FSSRDM .1858.1858 .5048.5048 .1390.1390 .7140.7140 .1855.1855 .8822.8822 .1076.1076 .3393.3393 .1962.1962 .7662.7662 .1973.1973 .7653.7653 .09114.09114 .4537.4537 .1501.1501 .7155.7155
SCFS .0951.0951 .2161.2161 .1387.1387 .6932.6932 .1859.1859 .8781.8781 .1077.1077 .3957.3957 .1978.1978 .7365.7365 .1944.1944 .7070.7070 .0928.0928 .4509.4509 .1447.1447 .7046.7046
LEDA .2053.2053 .5226.5226 .1401.1401 .7242.7242 .2044.2044 .9461.9461 .1065.1065 .3312.3312 .1944.1944 .7610.7610 .1951.1951 .7397.7397 .0915.0915 .4559.4559 .1488.1488 .7030.7030
LSMR .1966.1966 .5048.5048 .1139.1139 .6317.6317 .1857.1857 .8715.8715 .1070.1070 .3318.3318 .1910.1910 .7366.7366 .1962.1962 .7263.7263 .0922.0922 .4533.4533 .1473.1473 .7167.7167
SMLFS .1050.1050 .2516.2516 .1102.1102 .5792.5792 .1805.1805 .8585.8585 .1044.1044 .3220.3220 .2012.2012 .7662.7662 .2023.2023 .7417.7417 .0943.0943 .4559.4559 .1456.1456 .7123.7123
SFAM .1069.1069 .2258.2258 .0937.0937 .5031.5031 .1945.1945 .9027.9027 .1064.1064 .3673.3673 .1710.1710 .6356.6356 .1995.1995 .7332.7332 .0917.0917 .4451.4451 .1503.1503 .7414.7414
SMDR-IC .1317.1317 .3839.3839 .1382.1382 .7109.7109 .1781.1781 .8884.8884 .1104.1104 .3984.3984 .2054.2054 .7667.7667 .2033.2033 .7447.7447 .0955.0955 .4581.4581 .1541.1541 .7501.7501
SFS-BLL .0951.0951 .1952.1952 .0643.0643 .3249.3249 .2044.2044 .9457.9457 .1021.1021 .3105.3105 .1959.1959 .7632.7632 .1955.1955 .7319.7319 .0901.0901 .4579.4579 .1477.1477 .7042.7042
SMDR-MRC .1062.1062 .2339.2339 .0551.0551 .2628.2628 .1884.1884 .8843.8843 .1027.1027 .3430.3430 .1711.1711 .7231.7231 .1739.1739 .6278.6278 .0880.0880 .4472.4472 .1330.1330 .6609.6609
SMFS .0956.0956 .2097.2097 .0519.0519 .2741.2741 .1789.1789 .8795.8795 .1110.1110 .3975.3975 .1546.1546 .5794.5794 .1773.1773 .6237.6237 .0917.0917 .4404.4404 .1366.1366 .6669.6669
SFGR .0900.0900 .2258.2258 .0506.0506 .3007.3007 .1785.1785 .8779.8779 .1038.1038 .3328.3328 .1939.1939 .7328.7328 .2071.2071 .7640.7640 .0905.0905 .4473.4473 .1440.1440 .7393.7393
LFFS .1388.1388 .3436.3436 .1099.1099 .5669.5669 .1855.1855 .8822.8822 .1076.1076 .3393.3393 .1869.1869 .7060.7060 .1973.1973 .7353.7353 .0914.0914 .4537.4537 .1477.1477 .7051.7051
UAFS-BH .0956.0956 .2048.2048 .1256.1256 .6444.6444 .1739.1739 .8671.8671 .1085.1085 .3963.3963 .2073.2073 .7779.7779 .1862.1862 .6717.6717 .0914.0914 .4426.4426 .1361.1361 .6574.6574
6.2 Experimental Results and Performance Analysis
In this section, we demonstrate the superior performance of the proposed method Access-MFS compared with other competing methods in terms of AP, MaF, RL and OE. Table II summarizes the experimental results of various methods on eight multi-label datasets in terms of AP and MaF metrics, with the best performance highlighted in bold. From Table II, we can see that the proposed method Access-MFS consistently outperforms other methods across all datasets. As to VirusGO and Recreation datasets, Access-MFS achieves an average improvement of over 9% in both AP and MaF metrics compared to all other methods. On Medical dataset, Access-MFS gains over 26% and 12% average improvement in terms of AP and MaF, respectively. As to Slashdot dataset, Access-MFS outperforms other methods with more than 12% improvement in average AP and MaF. For Science dataset, Access-MFS achieves an average improvement of over 6% in AP and 4% in MaF, respectively. On the Computer dataset, Access-MFS yields average improvement exceeding 2.5% in AP and 5.7% in MaF. On Langlog and Enron datasets, Access-MFS achieves an average improvement of over 3.6% in AP and still outperforms all other methods in MaF. Additionally, Table III presents the experimental results of various methods in terms of RL and OE. From this table, we observe that the proposed Access-MFS consistently outperforms all other methods across all datasets. Furthermore, compared to the baseline method All-Fea, Access-MFS is superior across all datasets in terms of AP, MaF, RL, and OE, demonstrating the effectiveness of Access-MFS. Access-MFS also shows better performance than the supervised method LEFS on all datasets, indicating the effectiveness of the proposed semi-supervised learning strategy in enhancing feature selection performance.
Due to the difficulty in determining the optimal number of selected features, we also present the performance (AP, MaF, RL and OE) of different methods as the number of selected features varies. Figs. 2, 3, 4 and 5 respectively show the AP, MaF, RL and OE values for different numbers of selected features, with the labeled instance ratio fixed at 40%. As illustrated in these four figures, the proposed Access-MFS method outperforms other methods in most cases, with the number of selected features varying from 100 to 200 on each dataset. Furthermore, we also investigate the performance of different methods with varying ratios of labeled instances. Figs. 6, 7, 8 and 9 respectively show the AP, MaF, RL and OE of different methods when the ratio of labeled instances ranges from 10% to 50%. From these figures, we can see that our method achieves better performance in most cases. In addition, our method performs better as the ratio of labeled instances increases, indicating that our approach effectively leverages labeled information to guide the collaborative learning of sample and label similarity graphs, ultimately enhancing feature selection performance. The superior performance of the proposed method is attributed to the adaptive collaborative learning of similarity graphs from both the sample and label perspectives, which are embedded into a generalized regression model with an extended uncorrelated constraint and a consistency constraint between predicted and existing labels.
@@ 6.2 Experimental Results and Performance Analysis
Figure 4: RL of different methods with varying numbers of selected features.
Figure 5: OE of different methods with varying numbers of selected features.
@@
Figure 8: RL of different methods on eight datasets with varying percentages of labeled training data.
Figure 9: OE of different methods on eight datasets with varying percentages of labeled training data.
@@ 6.3 Parameter Sensitivity Experiment
Figure 11: Convergence curves of Access-MFS.
@@ 6.4 Convergence Experiment
The theoretical convergence of the proposed Access-MFS algorithm has been established in Section 5.1. This subsection provides experimental validation of its convergence and analyzes the speed of convergence. Fig. 11 illustrates the convergence curves of Access-MFS across eight datasets, showing the variation in the objective function value over the course of iterations. As illustrated in Fig. 11, the convergence curves steeply fall within a few iterations and approach stability almost within 15 iterations. The experimental results demonstrate that the proposed method can converge effectively.